Shift report 20030124 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Jan 24 2003 - 08:10:53 EST

  • Next message: Apache: "Shift report 20030124 08:00-16:00"
    ------------------------------------------------------------
    Date: Fri Jan 24 08:10:53 2003, Shift: 00:00-08:00
    Supervisor: AM, with: 
    Log book#13 pages 124 - 125
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam was for the most part of the shift.
    2. Runs 7609-7613 have been recorded, finishing up 300k trig.6 events at FS at 12 degrees, 1/8 field, A-polarity.
    MRS at 90 deg., and 350 current.
    3. DAQ control program crashed from time to time (IOServer errors).
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri Jan 24 2003 - 08:11:45 EST